OpenStudy (anonymous):

a gas has a pressure of 120 kpa, a temperature of 400 k and a volume of 50 milliliters. what volume will the gas have at a pressure of 60 kpa and a temperature of 200 kpa?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Gas_laws You want to use P1*V1/T1 = P2*V2/T2 120 * 50 / 400 = 60 * X / 200 -> solve for x I am assuming you meant 200K, not 200 Kpa

OpenStudy (anonymous):

Oh yeah sorry and thank you I'll try it out